OpenStudy (anonymous):

An equation is shown below: 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

well i am not so good with theory but i guess what are you supposed to do here is write this equation in different ways for 2nd part write it in terms of x and y

OpenStudy (anonymous):

lets see if someone else can be helpful @kirbykirby @Gabylovesyou

OpenStudy (anonymous):

inconsistent equations are those which cant be solved so this equation 9x - y = -2 multiply it by '-' so -(9x - y = -2) -9x+y=2 so you have two equations now 9x - y = -2 and -9x+y=2 both these equations cant be solved so they are inconsistent

OpenStudy (anonymous):

thats just one example multiply or divide it by anything and you will get inconsistent system of equations

OpenStudy (anonymous):

and i dont know what exactly the first question is for 2nd you do have to write it as x= .... and y = ... i guess

OpenStudy (anonymous):

@ganeshie8